The mean and standard deviation for the life time of a population of light bulbs are
1200 and 50 hours respectively. Assuming these lifetimes are normally distributed,
what is the probability that a light bulb will last over 1500 hours?
1
Expert's answer
2020-08-17T19:11:13-0400

"P(X>1500)=P(Z>\frac{1500-1200}{50})=P(Z>6)=1-P(Z<6)="

"=1-1=0."